OpenStudy (anonymous):

After 55.0 min, 34.0% of a compound has decomposed. What is the half-life of this reaction assuming first-order kinetics?

OpenStudy (abb0t):

Beginning amount = 100 Final amount = 100-41 Elapsed time = 55 min I believe if I have it right: \[half-life = \frac{ \log2(elapsed time) }{ \log(\frac{ {initial} }{ final }) }\] \[final-amount = \frac{ initial }{ 2^n }\] where \[n = \frac{ elapsed-time }{ half-life }\]